A ray of sunlight is passing from diamond into crown glass; theangle of incidence is 31.00°. Theindices of refraction for the blue and red components of the rayare: blue (ndiamond = 2.444,ncrown glass = 1.531), and red(ndiamond = 2.410, ncrownglass = 1.520). Determine the angle between the refractedblue and red rays in the crown glass.
Answer #1
Given that
A ray of sunlight is passing from diamond into crown glass;the angle of incidence is(θi) =31.00°
The indices of refraction for the blue and red components ofthe ray are
       Blue(ndiamond =2.444, ncrown glass = 1.531),
And Red (ndiamond =2.410, ncrown glass = 1.520).
Determine the angle between the refracted blue and red(θbluered) =?
Now when a ray of sunlight is passing from diamond into crownglass then the angle of refraction is given for the bluelight is
                           ndiamondsinθi =ncrownsinθr
Therefore the refracted angle for the blue light is
                           sinθr =   ndiamondsinθi /ncrown
                                θr=sin-1(ndiamondsinθi/ncrown)
                                   = sin-1((2.444)sin(31.00°)/1.531)
                                   = sin-1(0.8221)
                                   = 55.30330
Now when a ray ofsunlight is passing from diamond into crown glass then the angle ofrefraction is given for the red lightis
                           ndiamondsinθi =ncrownsinθr Therefore the refractedangle for the red light is
                           sinθr =   ndiamondsinθi/ncrown
                                θr=sin-1(ndiamondsinθi/ncrown)
                                   = sin-1((2.410)sin(31.00°)/1.520)
                                   = sin-1(0.8166)
                                   = 54.74650
The angle between the refracted blue and red(θbluered) =55.30330 - 54.74650
                                                                                                                     = 0.55680